OpenStudy (anonymous):

Solve by elimination I don't need the answer I need the work -2x+2y+3z=0 -2x-y+z=-3 2x+3y+3z=5

OpenStudy (campbell_st):

ok... so you will need to reduce this to 2 equations with 2 unknowns. it is probably easiest to eliminate x multiply every term in the 1st equation by -1 then add it to the 2nd equation Call this equation 4 add the 2nd and 3rd equations...again eliminating x and call it equation 5 now you have 2 equations in 2 unknowns... and you need to select a new pronumeral to eliminate

OpenStudy (anonymous):

I think my brain just died. but lets try this

OpenStudy (anonymous):

2y x -1=-2y 3z x -1=-3z 0 x -1=0

OpenStudy (campbell_st):

here is an easier way add equation 1 and 3 -2x + 2y + 3z = 0 2x + 3y + 3z = 5 ------------------ 5y + 6z = 5 call it equation 4 next add equation 2 and equation 3 -2x - y + z = -3 2x + 3y + 3z = 5 -------------------- 2y + 4z = 2 halve this equation y + 2z = 1 call it equation 5 now eliminate z equation 4 + (-3) times equation 5 5y + 6z = 5 -3y -6z = --3 ------------------------- 2y = 2 now you can find y. to find z substitute into either equation 4 or 5 to find x, substitute into any of the original equations.